OpenStudy (anonymous):

Solve the following equations for x. Identify any extraneous solutions. Show all work.-4√(x+9)=20 Pleaseeeeee

OpenStudy (zale101):

-4√(x + 9) = 20 √(x + 9) = -5 so what do you think, does it have a solution or not

OpenStudy (anonymous):

im going to say yes but im really confused i dont know how to even start this problem

OpenStudy (anonymous):

Does x=16?

OpenStudy (anonymous):

im not sure. how did you get that?

OpenStudy (zale101):

√(x+9) = -5 x+9 = 25 x = 16

OpenStudy (anonymous):

Is that it? That seems kinda easy. :)

OpenStudy (anonymous):

You get rid of the square root by squaring both sides

OpenStudy (zale101):

So, what do you think. Does it have a solution ?

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

yes

OpenStudy (zale101):

there's no solution

OpenStudy (anonymous):

Hmm. square root of 25 equals both positive and negative 5...

OpenStudy (anonymous):

Thanks for all the help Zale and Jenny :)

OpenStudy (anonymous):

Np

OpenStudy (zale101):

you're welcome :)